# Kudzu: Fast and Simple High-Throughput BFT

### v1

@@ 5. Protocol Analysis
By the Quorum Size Property (see Section 3.3.1) for notarization certificates, at least $n-2f-p$ honest replicas must have broadcast notarization shares for $B$. Since we are assuming $n\geq 3f+2p+1$, it follows that $n-2f-p>0$, so some honest replica $P$ must have broadcast a notarization share for $B$. This could happen either at 22 or at line 36. In the first case, $B$ must be the block that $P$ received as a proposal from the leader. In the second case, since $P$ received $f+p+1$ first votes for $B$, one of these must be a first vote for $B$ from some honest replica $Q$, and so $B$ must be the block that $Q$ received as a proposal from the leader. ∎
@@
We are relying on the Quorum Size Property (see Section 3.3.1) for notarization certificates: when a notarization certificate for a block $B$ is added to the certificate pool, at least $n-2f-p$ honest replicas must have already broadcast notarization votes for $B$, which contain $B$ as well as fragments sufficient to reconstruct $B$’s payload, since $n-2f-p\geq f+p+1$. ∎
@@ 5.1. Safety
If an honest replica receives $f+p+1$ first votes for a block $B$ in round $v$, then no block different from $B$ can be fast finalized in round $v$ by any honest replica.
@@
If the inequality $\textsf{allVotes}(\textit{firstVote})-\textsf{maxVotes}(\textit{firstVote})\geq f+p+1$ holds for a replica in round $v$, then no block can be fast finalized in round $v$ by any replica.
@@
We first prove the Incompatibility of Notarization and Fast Finalization. Suppose towards contradiction that for slot $v$ a fast finalization certificate exists for block $B$ and a notarization certificate exists for block $C$, with $C\neq B$. By the Quorum Size Property (see Section 3.3.1) for notarization certificates, this implies that an honest replica broadcast a notarization vote for $C$.
  • On the one hand, suppose that the notarization vote for $C$ was sent by the protocol on 36. Due to the condition on 33, this means that $f+p+1$ first votes were received for a block $D$. Note that $D$ is a non-timeout block. Moreover, by the logic of ReconstructAndNotarize, either $C$ is a timeout block or $D=C$. By the Uniqueness of Fast Finalization Property (Lemma 5.5) we know that $D=B$. Since $B$ is assumed to be valid, $C$ cannot be a timeout block, so we also have $B=D=C$, a contradiction.
  • On the other hand, suppose that the notarization vote for $C$ was sent on 40. By the Absence of Fast Finalization Property (Lemma 5.6), this implies that no block is fast finalized, again a contradiction.
The Incompatibility of Notarization and Finalization Property follows from a standard quorum intersection argument, based on the fact that in each slot an honest replica issues a finalization vote only for a block only if it did not send a notarization vote for a different block in that slot (see 10). Suppose towards contradiction that for slot $v$ a finalization certificate exists for block $B$ and a notarization certificate exists for block $C$, with $C\neq B$. By the Quorum Size property (see Section 3.3.1) for finalization and notarization certificates, if $f^{\prime}\leq f$ is the number of corrupt replicas, then at least $n-f-p-f^{\prime}$ honest replicas broadcast finalization votes for $B$, and a disjoint set of at least the same number of honest replicas broadcast notarization votes for $C$. This implies that there are at least $2(n-f-p-f^{\prime})$ distinct honest replicas. However, under the assumption that $n\geq 3f+2p+1$, we have $12(n-f-p-f^{\prime})\geq n-f^{\prime}+$ 1, a contradiction. ∎
@@ 5.2. Liveness
Consider a slot $v\geq 1$ and suppose the leader for slot $v$ is an honest replica $Q$. Suppose that the first honest replica $P$ to enter the loop iteration for slot $v$ does so at time $T_{0}$. Suppose that the network is $\delta$-synchronous over the interval $[T_{0},T_{0}+4\delta]$ for some $\delta$ with $\Delta_{\text{\rm timeout}}\geq 2\delta$. Then, $Q$ will propose a block for slot $v$ by time $T\leq T_{0}+\delta$. Each honest replica will finish the loop iteration before time $T+2\delta$ by adding $Q$’s proposed block $B$ to its complete block tree. Moreover, if $n-p$ replicas are honest, each honest replica will finalize $B$ by time $T+2\delta$. If more than $p$ replicas are corrupt, each honest replica will finalize $B$ by time $T+3\delta$.
@@
By the Completeness Properties (Lemma 5.2 and Lemma 5.3), before time $T_{0}+\delta$, each honest replica will enter slot $v$ by time $T_{0}+\delta$, having either a timeout certificate for slot $v-1$ or a block for slot $v-1$ in its complete block tree. Before time $T\leq T_{0}+\delta$, the leader $Q$ will propose a block $B$ that extends a block $B^{\prime}$ with slot number $v^{\prime}<v$. By the logic of the protocol, we know that $Q$ must have timeout certificates for slots $v^{\prime}+1,\ldots,v-1$ at the time it makes its proposal, as well as a notarization certificate for $B^{\prime}$. Again by the Completeness Properties, before time $T+\delta$, each honest replica will have $B^{\prime}$ in its complete block tree and all of these timeout certificates in its certificate pool. Each honest replica will receive $B$ before this time, and because $\Delta_{\text{\rm timeout}}\geq 2\delta$, will broadcast a first vote for $B$ by this time. Because all honest replicas broadcast a first vote for $B$, each such replica will only see ever see at most $f$ first votes for any other block. It follows that each honest replica will only ever see
@@
[…]
Therefore, honest replicas will not broadcast a notarization votes in slot $v$ for anything other than $B$. Before time $T+2\delta$, each honest replica will have added $B$ to its complete block tree and broadcast a finalization vote on $B$. If $n-p$ replicas are honest, each honest replica will have added a fast finalization certificate to its pool by time $T+2\delta$ as well. Otherwise, if more than $p$ replicas are corrupt, each honest replica will finalize $B$ before time $T+3\delta$, when adding the finalization certificate for slot $v$ to its pool. ∎
The second lemma analyzes the pessimistic case, when the network is asynchronous or the leader of a given round is corrupt. It says that eventually, all honest replicas will move on to the next round.
@@
Suppose that the network is $\delta$-synchronous over an interval $[T,T+\Delta_{\text{\rm timeout}}+3\delta]$, for an arbitrary value of $\delta$, and that at time $T$, some honest replica is in the loop iteration for slot $v$ and all other honest replicas are in a loop iteration for $v$ or a previous slot. Then, before time $T+\Delta_{\text{\rm timeout}}+3\delta$, all honest replicas exit slot $v$.
@@
[…]
(b). Consider some honest replica $P$. By time $T+\Delta_{\text{\rm timeout}}+2\delta$, $P$ will have observed all votes of other honest replicas. If the only entries in firstVote are those from honest replicas, then the inequality $\textsf{allVotes}(\textit{firstVote})-\textsf{maxVotes}(\textit{firstVote})\geq f+p+1$ must hold. To see this, if $f^{\prime}\leq f$ is the number of corrupt replicas, then
@@
$\geq f+p+1\quad\text{(since $n\geq 3f+2p+1$).}$.
@@ 5.3. Boundedness
[…]
@@
So it suffices to show that $(3(n-f))/(n-2f-p)<6$. This is easily seen to follow by a simple calculation using (1).
@@
This immediately gives us bounds on the message and storage complexity if the protocol per slot.
@@ 5.4. Complexity
Based on the concrete bounds in Section 5.3 (and the preliminary discussion in Section 4.2), it is easily seen that the message complexity per slot is $O(n^{2})$. Based on the properties of erasure codes and Merkle trees discussed in Section 3.3.2, the communication complexity per slot is $O(\beta n+n^{2}\log(n)\kappa+n^{2}\lambda)$, where $\beta$ is a bound on the payload size, $\kappa$ is the output length of the collision-resistant hash, and $\lambda$ is a bound on the length of any signature share or certificate. Moreover, the communication is balanced, in that every replica, including the leader, transmits the same amount of data, up to a constant facrtor. It is also easily seen that each replica needs to store $O(\beta+n\log(n)\kappa+n\lambda)$ bits of data.
@@ 6. Protocol Variations
Our protocol can be adapted to the setting of $n\geq 3f+2p^{*}-1$, where $p^{*}\geq 1$  (abraham2021good, ), with the insight that if honest replicas vote for different blocks in the same slot, the leader has to be corrupt. The liveness analysis can leverage the fact that, in this case, honest replicas will observe at least $n-f$ votes from non-leader replicas, as is done in (banyan, ). We plan to analyze a variation of Kudzu with this adaptation in an extended version of this paper.
A variation of DispersedSimplex (song, ) features segments of consecutive slots with the same leader. Such stable leader assignment is beneficial for the throughput of the protocol. The same technique can be applied to our protocol, and we plan to analyze a variation of Kudzu featuring stable leaders in an extended version of this paper.

### v2

@@ 5. Protocol Analysis
By the Quorum Size Property (see Section 3.3.1) for notarization certificates, at least n2fpn-2f-p honest replicas must have broadcast notarization shares for $B$. Since we are assuming $n\geq 3f+2p+1$, it follows that n2fp>0n-2f-p>0, so some honest replica $P$ must have broadcast a notarization share for $B$. This could happen either at line 24 or at line 38. In the first case, $B$ must be the block that $P$ received as a proposal from the leader. In the second case, since $P$ received $f+p+1$ first votes for $B$, one of these must be a first vote for $B$ from some honest replica $Q$, and so $B$ must be the block that $Q$ received as a proposal from the leader. ∎
@@
We are relying on the Quorum Size Property (see Section 3.3.1) for notarization certificates: when a notarization certificate for a block $B$ is added to the certificate pool, at least n2fpn-2f-p honest replicas must have already broadcast notarization votes for $B$, which contain $B$ as well as fragments sufficient to reconstruct $B$’s payload, since $n-2f-p\geq f+p+1$. ∎
@@ 5.1. Safety
If an honest replica receives $f+p+1$ first votes for a block $B$ in slot $v$, then no block different from $B$ can be fast finalized in slot $v$ by any honest replica.
@@
If the inequality $\textsf{allVotes}(\textit{firstVote})-\textsf{maxVotes}(\textit{firstVote})\geq f+p+1$ holds for a replica in slot $v$, then no block can be fast finalized in slot $v$ by any replica.
@@
We first prove the Incompatibility of Notarization and Fast Finalization. Suppose towards contradiction that for slot $v$ a fast finalization certificate exists for block $B$ and a notarization certificate exists for block $C$, with $C\neq B$. By Lemma 5.4 at most $p$ honest replicas first vote for anything other than $B$. By the Quorum Size Property (see Section 3.3.1) at least $(n-f-p)-f\geq f+p+1$ honest replicas must send a notarization vote for $C$. Thus, at least $f+1$ honest replicas must send a notarization vote for $C$ after having sent a first vote.
  • On the one hand, suppose that a notarization vote for $C$ was sent by the protocol on line 38. Due to the condition on line 35, this means that $f+p+1$ first votes were received for a block $D$. Note that $D$ is a non-timeout block. Moreover, by the logic of ReconstructAndNotarize, either $C$ is a timeout block or $D=C$. By the Uniqueness of Fast Finalization Property (Lemma 5.5) we know that $D=B$. Since $B$ is assumed to be valid, $C$ cannot be a timeout block, so we also have $B=D=C$, a contradiction.
  • On the other hand, suppose that a notarization vote for $C$ was sent on line 42. By the Absence of Fast Finalization Property (Lemma 5.6), this implies that no block is fast finalized, again a contradiction.
The Incompatibility of Notarization and Finalization Property follows from a standard quorum intersection argument, based on the fact that in each slot an honest replica issues a finalization vote only for a block only if it did not send a notarization vote for a different block in that slot (see line 12). Suppose towards contradiction that for slot $v$ a finalization certificate exists for block $B$ and a notarization certificate exists for block $C$, with $C\neq B$. By the Quorum Size property (see Section 3.3.1) for finalization and notarization certificates, if $f^{\prime}\leq f$ is the number of corrupt replicas, then at least $n-f-p-f^{\prime}$ honest replicas broadcast finalization votes for $B$, and a disjoint set of at least the same number of honest replicas broadcast notarization votes for $C$. This implies that there are at least $2(n-f-p-f^{\prime})$ distinct honest replicas. However, under the assumption that $n\geq 3f+2p+1$, we have $2(n-f-p-f^{\prime})\geq n-f^{\prime}+1$, a contradiction. ∎
@@ 5.2. Liveness
Consider a slot $v\geq 1$ and suppose the leader for slot $v$ is an honest replica $Q$. Suppose that the first honest replica $P$ to enter the loop iteration for slot $v$ does so at time $T_{0}$. Suppose that the network is $\delta$-synchronous over the interval $T_{0},T_{0}+4\delta]$ for some $\delta$ with $\Delta_{\text{\rm timeout}}\geq 2\delta$. Then, $Q$ will propose a block for slot $v$ by time $T\leq T_{0}+\delta$. Each honest replica will finish the loop iteration before time $T+2\delta$ by adding $Q$’s proposed block $B$ to its complete block tree. Moreover, if npn-p replicas are honest, each honest replica will finalize $B$ by time $T+2\delta$. If more than $p$ replicas are corrupt, each honest replica will finalize $B$ by time $T+3\delta$.
@@
By the Completeness Properties (Lemma 5.2 and Lemma 5.3), before time $T_{0}+\delta$, each honest replica will enter slot $v$ by time $T_{0}+\delta$, having either a timeout certificate for slot v1v-1 or a block for slot v1v-1 in its complete block tree. Before time $T\leq T_{0}+\delta$, the leader $Q$ will propose a block $B$ that extends a block $B^{\prime}$ with slot number $v^{\prime}<v$. By the logic of the protocol, we know that $Q$ must have timeout certificates for slots $v^{\prime}+1,\ldots,v-1$ at the time it makes its proposal, as well as a notarization certificate for $B^{\prime}$. Again by the Completeness Properties, before time $T+\delta$, each honest replica will have $B^{\prime}$ in its complete block tree and all of these timeout certificates in its certificate pool. Each honest replica will receive $B$ before this time, and because $\Delta_{\text{\rm timeout}}\geq 2\delta$, will broadcast a first vote for $B$ by this time. Because all honest replicas broadcast a first vote for $B$, each such replica will only see ever see at most $f$ first votes for any other block. It follows that each honest replica will only ever see
@@
  • $\textsf{allVotes}(\textit{firstVote})-\textsf{maxVotes}(\textit{firstVote})\leq\textsf{allVotes}(\textit{firstVote})-\mathit{count}_{B}\leq f,$ where $\mathit{count}_{B}$ is the number of of first votes for $B$ that it sees.
Therefore, honest replicas will not broadcast a notarization votes in slot $v$ for anything other than $B$. Before time $T+2\delta$, each honest replica will have added $B$ to its complete block tree and broadcast a finalization vote on $B$. If npn-p replicas are honest, each honest replica will have added a fast finalization certificate to its pool by time $T+2\delta$ as well. Otherwise, if more than $p$ replicas are corrupt, each honest replica will finalize $B$ before time $T+3\delta$, when adding the finalization certificate for slot $v$ to its pool. ∎
The second lemma analyzes the pessimistic case, when the network is asynchronous or the leader of a given slot is corrupt. It says that eventually, all honest replicas will move on to the next slot.
@@
Suppose that the network is $\delta$-synchronous over an interval $T,T+\Delta_{\text{\rm timeout}}+3\delta]$, for an arbitrary value of $\delta$, and that at time $T$, some honest replica is in the loop iteration for slot $v$ and all other honest replicas are in a loop iteration for $v$ or a previous slot. Then, before time $T+\Delta_{\text{\rm timeout}}+3\delta$, all honest replicas exit slot $v$.
@@
Case (a). Since $f+p+1$ replicas cast notarization votes for $B$, some honest replica did so. Since this replica had to have $B$’s parent in its complete block tree, by Completeness Properties (Lemma 5.3) each honest replica will have $B$’s parent in its complete block tree by time $T+\Delta_{\text{\rm timeout}}+2\delta$. All honest replicas observe all first votes from other honest replicas before time $T+\Delta_{\text{\rm timeout}}+2\delta$, and so each honest replica will call $\textsf{ReconstructAndNotarize}(v,B)$ before that time, unless it has already exited slot $v$. If some honest replica has exited slot $v$ before that time, the by Completeness Properties (Lemma 5.2 and Lemma 5.3), all honest replicas will exit the slot before time $T+\Delta_{\text{\rm timeout}}+3\delta$. Otherwise, assume no honest replica has exited before time $T+\Delta_{\text{\rm timeout}}+2\delta$. Whenever $\textsf{ReconstructAndNotarize}(v,B)$ is called by some honest replica, it has received at least $f+p+1$ first votes for $B$, and so can attempt to reconstruct $B$. If it fails to reconstruct $B$’s payload or finds that it is invalid, then it and all honest replicas will do so and issue a timeout vote (this follows from collision resistance of the hash function and the fragment decoding logic). Otherwise, each honest replica will issue a notarization vote for $B$. Therefore, before time $T+\Delta_{\text{\rm timeout}}+3\delta$, each honest replica will either add $B$ to its complete block tree or the timeout certificate to its pool, and proceed exit the slot.
Case (b). Consider some honest replica $P$. By time $T+\Delta_{\text{\rm timeout}}+2\delta$, $P$ will have observed all votes of other honest replicas. If the only entries in firstVote are those from honest replicas, then the inequality $\textsf{allVotes}(\textit{firstVote})-\textsf{maxVotes}(\textit{firstVote})\geq f+p+1$ must hold. To see this, if $f^{\prime}\leq f$ is the number of corrupt replicas, then
@@
$\geq f+p+1\quad\text{(since $n\geq 3f+2p+1$).}$
@@ 5.3. Boundedness
Let us consider notarization votes on non-timeout blocks. An honest replica sends a first vote for at most one such block. Any notarization vote for some other non-timeout block $B$ requires that the replica found $B\in\textsf{manyVotes}(\mathit{firstVotes})$. In other words, the replica has seen at least $f+p+1$ other replica’s first votes for $B$. At most one first vote per replica is considered when computing $\textsf{manyVotes}(\mathit{firstVotes})$. Therefore, an honest replica can cast at most $\lfloor n/(f+p+1)\rfloor$ notarization votes beyond the first vote, and by (2), $\lfloor n/(f+p+1)\rfloor\leq 2$, for a total of 3 notarization votes for non-timeout blocks.
@@
So it suffices to show that (3(nf))/(n2fp)<6(3(n-f))/(n-2f-p)<6. This is easily seen to follow by a simple calculation using (1).
@@
This immediately gives us bounds on the message and storage complexity of the protocol per slot.
@@ 5.4. Complexity
Based on the concrete bounds in Section 5.3 (and the preliminary discussion in Section 4.2), it is easily seen that the message complexity per slot is $O(n^{2})$. Based on the properties of erasure codes and Merkle trees discussed in Section 3.3.2, the communication complexity per slot is $O(\beta n+n^{2}\log(n)\kappa+n^{2}\lambda)$, where $\beta$ is a bound on the payload size, $\kappa$ is the output length of the collision-resistant hash, and $\lambda$ is a bound on the length of any signature share or certificate. Moreover, the communication is balanced, in that every replica, including the leader, transmits the same amount of data, up to a constant factor. It is also easily seen that each replica needs to store $O(\beta+n\log(n)\kappa+n\lambda)$ bits of data.
@@ 6. Protocol Variations
Our protocol can be adapted to the setting of $n\geq 3f+2p^{*}-1$, where $p^{*}\geq 1$  (abraham2021good, ), with the insight that if honest replicas vote for different blocks in the same slot, the leader has to be corrupt. The liveness analysis can leverage the fact that, in this case, honest replicas will observe at least nfn-f votes from non-leader replicas, as is done in (banyan, ). We plan to analyze a variation of Kudzu with this adaptation in follow-up work.
A variation of DispersedSimplex (song, ) features segments of consecutive slots with the same leader. Such stable leader assignment is beneficial for the throughput of the protocol. The same technique can be applied to our protocol, and we plan to analyze a variation of Kudzu featuring stable leaders in follow-up work.